OpenStudy (anonymous):

Can someone please help, I have been stuck on this for a while. Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). y = one thirdx − 3 y = one thirdx + 17 y = −3x − 3 y = −3x + 17

OpenStudy (anonymous):

Remember the slope of the perpendicular lines: \[(\frac{ c }{ d })(\frac{ -d}{ c }) = -1\]

OpenStudy (anonymous):

Ok well the formula you want is: \[y-y _{1} = m(x-x_{1})\] So you have: \[3x+y=7\]\[y= -3x + 7\] So slope would be the negative reciprocal of -3, which is: \[\frac{ 1 }{ 3 }\] Your equation should look like: \[y-(-1)=\frac{ 1 }{ 3 }(x-6)\] Should be easy from here.

OpenStudy (anonymous):

Ok, I am going to do the work, sorry it took so long for me to see this.

OpenStudy (anonymous):

Is it y= -3x + 17?